### v3

@@ 5 Discussion
The findings from this systematic review reveal several key trends and challenges in the detection of SATD over the past decade, demonstrating both the evolution and fragmentation within the field. SATD detection approaches have progressed significantly, shifting from simple keyword-based NLP methods to more complex ML, DL, and Transformers-based models. While each advancement has brought increased performance and automation, the development of practical and scalable SATD detection tools remains limited.
5.1 Evolution of SATD detection techniques
Early SATD detection techniques primarily relied on basic NLP methods, such as keyword matching and textual patterns, to identify technical debt in code comments. These approaches, although foundational, were often limited by their inability to account for the nuanced language used by developers and their reliance on rule-based systems that could not easily adapt to different contexts. From 2017 onward, ML techniques such as Naive Bayes (NB), Support Vector Machines (SVM), and Random Forests became more prevalent, offering improved accuracy by learning from labeled datasets and reducing false positives. Despite this progress, these methods still faced challenges related to generalizability across different datasets and projects, as well as difficulty in explaining the rationale behind the classification of certain comments as SATD.
The rise of deep learning methods from 2020 to 2022 marked a significant leap forward in SATD detection. Models like Long Short-Term Memory (LSTM) networks and Convolutional Neural Networks (CNN) enabled the identification of more complex patterns in developer comments and project artifacts. These approaches improved recall and reduced false negatives, offering a more balanced approach to identifying SATD across diverse datasets. However, the high computational cost of training and deploying these models remains a major barrier to widespread adoption.
Most recently, Transformers-based models, such as BERT and RoBERTa, have emerged as the leading approaches for SATD detection. These models have demonstrated superior performance in capturing the context and semantics of developer comments, enabling more accurate and generalizable SATD identification. The ability of Transformers to handle large-scale datasets and complex text has pushed the boundaries of what is possible in SATD detection. However, issues related to the explainability of these models and their scalability for industrial use still persist.
@@ 5.2 Challenges in SATD detection
Despite the advancements in SATD detection techniques, several challenges hinder the practical adoption of these methods in real-world software development environments. A key challenge in SATD detection is the imbalance in datasets—SATD instances are often much less frequent than non-SATD instances, which negatively impacts model performance. This imbalance is not limited to the distinction between SATD and non-SATD but also occurs among specific SATD types, where certain categories, such as requirements debt or design debt, are far less represented, further complicating detection. Although several techniques have been employed to address data imbalance in SATD detection, future research should explore more advanced methods, such as leveraging large language models (LLMs) to better handle imbalanced data and improve model performance.
Generalizability remains another key challenge. Many SATD detection models are trained on specific datasets, often from open-source projects, which limits their ability to perform well on new or proprietary codebases. This lack of generalizability reduces the practical utility of these models in diverse software development environments, where coding practices, project documentation, and domain-specific language can vary significantly. The explainability of advanced models, particularly deep learning and Transformer-based approaches, is also a concern. Without a clear understanding of why certain comments are classified as SATD, developers may be reluctant to trust and adopt these models in their workflows. Future work should focus on developing explainable AI techniques for SATD detection to ensure that these models’ decision-making processes are transparent and understandable to developers. Research on enhancing the interpretability of DL and Transformers models, perhaps through visualizations or simplified intermediate representations, would help practitioners better trust and utilize these models in real-world software development settings.
Another significant barrier to adoption is the high computational cost associated with training and maintaining deep learning and Transformers-based models. While these models offer improved accuracy, their resource-intensive nature makes them difficult to implement for smaller development teams or in environments with limited computational resources. This challenge highlights the need for more efficient algorithms or alternative approaches that can deliver high accuracy without the associated computational burden.
@@ 5.3 Lack of tools for SATD detection
Despite the significant academic interest in SATD detection, the translation of research into practical tools remains limited. This review identified only seven tools developed specifically for SATD detection over the past decade, reflecting a notable gap between theoretical research and practical application. While tools like SATD Detector and eXcomment provide valuable functionality for identifying SATD in Java codebases, the overall number of available tools is insufficient to meet the needs of diverse software development teams. Additionally, the tools that do exist often lack the ability to integrate seamlessly into modern development environments, limiting their utility for real-time SATD management.
The limited development of practical tools also highlights the need for further collaboration between academia and industry. Researchers must focus on creating tools that not only demonstrate high performance in academic settings but also address the specific needs and constraints of industrial software development, such as scalability, usability, and integration with existing development environments. Bridging the gap between theoretical advancements and practical, scalable tools is essential for achieving widespread adoption in industry. Future research should prioritize developing user-friendly tools that can be integrated into continuous integration (CI) environments, enabling real-time SATD detection and providing actionable insights for developers. Moreover, these tools should facilitate collaboration within teams working on large-scale software projects, ensuring that they are both practical and effective in real-world development scenarios.
@@ 5.4 Future Directions
To address the challenges identified in this review, future research should prioritize enhancing the generalizability, explainability, and efficiency of SATD detection models. Techniques such as transfer learning, which allow models to be trained on one dataset and then applied to others with minimal retraining, could significantly improve generalizability. At the same time, incorporating explainable AI (XAI) techniques is crucial to ensure that the decision-making processes of these models are transparent and understandable for developers. Developing more interpretable models, or employing post-hoc explanation techniques, could foster trust in these tools and encourage their adoption in real-world projects. Further research on enhancing the interpretability of deep learning and Transformer models—potentially through visualizations or simplified intermediate representations—would empower practitioners to better understand, trust, and effectively utilize these models in software development settings.
There is also a need for the development of more practical SATD detection tools that can be easily integrated into existing development environments. These tools should focus on providing real-time feedback to developers, enabling them to address technical debt as part of their normal workflow. By bridging the gap between research and practice, the software engineering community can ensure that SATD detection techniques contribute meaningfully to the long-term maintenance and evolution of software systems.
@@ 6 Threats to Validity
As with any systematic literature review (SLR), the results of this study are subject to various validity threats. These are categorized as internal, external, and construct validity threats, along with mitigation strategies to minimize their impact.
@@ 6.1 Internal validity
Internal validity refers to the reliability of the methodology employed in the study. One key threat to internal validity in this SLR is related to the formulation of the search strings. In the pilot search, several iterations of search strings were tested, starting with a more general search (e.g., “self-admitted technical debt" OR SATD and NLP), which led to a broader, less focused set of results. Over time, more refined strings were developed, including terms like “detection" and “identification," as well as the use of boolean operators and the wildcard (*) for broader term variations.
The final search string (S3), which combined terms such as “self-admitted technical debt" OR SATD with NLP OR “natural language processing" AND (detection OR identification OR prediction), was chosen after multiple pilot tests to ensure it captured the most relevant studies. This iterative process helped mitigate the risk of missing key studies during the initial literature search.
@@ 6.2 External validity
External validity refers to the extent to which the findings of this SLR can be generalized to other contexts beyond the studies reviewed. A primary threat to external validity in this research is the limited generalizability of the results due to the scope of the reviewed datasets. All of the studies included in this review rely on open-source projects, which may not fully capture the complexities or constraints faced in proprietary or large-scale industrial environments.
Open-source datasets often have characteristics, such as well-documented codebases or community-driven development practices, that differ significantly from proprietary software projects, which may involve stricter deadlines, unique domain requirements, or higher security and compliance constraints. As a result, some of the trends and findings related to SATD detection may not extend to these more complex and regulated environments. For example, the performance of machine learning or deep learning models, which may perform well on open-source datasets, could degrade when applied to more diverse or proprietary datasets with varying development practices.
@@ 6.3 Construct validity
Construct validity pertains to whether the study accurately measures the concepts it aims to investigate. A significant threat in this SLR is the comparison of results from different studies, which can be misleading if the studies use different methodologies or report metrics inconsistently. Directly comparing the precision, recall, and F1-scores of models trained on heterogeneous datasets without considering their context could lead to inaccurate conclusion.
To address this, we employed a grouping strategy to compare similar studies and ensured that the datasets were organized before any comparisons were made. By contextualizing the results within groups, we aimed to make the comparisons more meaningful. Furthermore, this methodology is based on established practices within the field, notably exemplified by Hall et al. 103 and Malhotra 104, which similarly undertakes the task of comparing results across different studies to deduce broad-based insights into algorithmic performance in varying contexts.

### v4

@@ 5 Discussion
The findings from this systematic review reveal several key trends and challenges in detecting SATD over the past decade, demonstrating both the evolution and fragmentation within the field. SATD detection techniques have progressed significantly, shifting from simple heuristic-based methods to more complex ML, DL, and Transformer-based models. While each advancement has brought increased performance and automation, the development of practical and scalable SATD detection tools remains limited.
5.1 Evolution of SATD detection techniques and artifacts
Early SATD detection techniques primarily relied on heuristic-based methods, such as keyword matching and textual patterns, to identify technical debt in code comments. These techniques, although foundational, were often limited by their inability to account for the nuanced language used by developers and their reliance on rule-based systems that could not easily adapt to different contexts. From 2017 onward, ML techniques such as NB, SVM, and RF became more prevalent, offering improved accuracy by learning from labeled datasets and reducing false positives. Despite this progress, these methods still faced challenges related to generalizability across different datasets and projects, as well as difficulty in explaining the rationale behind the classification of certain comments as SATD.
The rise of DL methods from 2020 to 2022 marked a significant leap forward in SATD detection. Models like LSTM and CNN enabled the identification of more complex patterns in developer comments and project artifacts. These techniques improved recall and reduced false negatives, thereby providing a more balanced technique for identifying SATD across diverse datasets. However, the high computational cost of training and deploying these models remains a major barrier to widespread adoption.
Most recently, Transformer-based models, such as BERT and RoBERTa, have emerged as the leading techniques for SATD detection. These models have demonstrated superior performance in capturing the context and semantics of developer comments, enabling more accurate and generalizable SATD identification. The ability of Transformer to handle large-scale datasets and complex text has pushed the boundaries of what is possible in SATD detection. However, issues related to the explainability of these models and their scalability for industrial use still persist.
It is also important to note that the evolution of these techniques has occurred primarily within the context of code comments, which remain the dominant artifact for SATD detection. While this focus is understandable given the accessibility and ubiquity of comments across software projects, it narrows the diversity and generalizability of existing approaches. Other artifacts, such as issue trackers, commit messages, and pull requests, have been comparatively underexplored despite their potential to provide complementary perspectives on technical debt. Addressing this imbalance represents a valuable direction for future research, as expanding the scope of artifacts can enable a more comprehensive understanding of SATD across the software development process.
@@ 5.2 Challenges in SATD detection
Despite the advancements in SATD detection techniques, several challenges hinder the practical adoption of these methods in real-world software development environments. A key challenge in SATD detection is the imbalance in datasets—SATD instances are often much less frequent than non-SATD instances, which negatively impacts model performance. This imbalance is not limited to the distinction between SATD and non-SATD but also occurs among specific SATD types, where certain categories, such as requirements debt or design debt, being far less represented, further complicating detection. Although several techniques have been employed to address data imbalance in SATD detection, future research should explore more advanced methods, such as leveraging large language models (LLMs) to better handle imbalanced data and improve model performance.
Generalizability remains another key challenge. Many SATD detection models are trained on specific datasets, often from open-source projects, which limits their performance on new or proprietary codebases. This lack of generalizability reduces the practical utility of these models in diverse software development environments, where coding practices, project documentation, and domain-specific language can vary significantly.
To address this, future research should move beyond within-project validation and systematically evaluate SATD detection models “in the wild”, using out-of-distribution datasets and cross-project settings. Such evaluations, as demonstrated by recent work like 109, are essential for understanding how well these models transfer to unseen contexts and for ensuring their reliability in real-world adoption.
The explainability of advanced models, particularly DL and Transformer-based techniques, is also a concern. Without a clear understanding of why certain texts are classified as SATD, developers may be reluctant to trust and adopt these models in their workflows. Future work should focus on developing explainable AI techniques for SATD detection to ensure that these models’ decision-making processes are transparent and understandable to developers. Research on enhancing the interpretability of DL and Transformer models, perhaps through visualizations or simplified intermediate representations, would help practitioners better trust and utilize these models in real-world software development settings.
Another significant barrier to adoption is the high computational cost associated with training and maintaining DL and Transformer-based models. While these models offer improved accuracy, their resource-intensive nature makes them difficult to implement for smaller development teams or in environments with limited computational resources. This challenge highlights the need for more efficient algorithms or alternative techniques to deliver high accuracy without the associated computational burden.
@@ 5.3 Lack of tools for SATD detection
Despite the significant academic interest in SATD detection, the translation of research into practical tools remains limited. Over the past decade, this systematic review identified only seven tools developed specifically for SATD detection. This reflects a notable gap between theoretical research and practical application and a lack of sufficient tools to meet the diverse needs of software development teams. While tools such as SATD Detector and eXcomment provide valuable functionality for identifying SATD in Java codebases, their scope is limited in terms of flexibility, scalability, and integration with modern software development workflows. Consequently, these tools have experienced limited adoption beyond academic prototypes and small-scale studies. The current state of SATD detection tools presents several challenges that hinder their widespread adoption among practitioners. Key areas of concern include limited real-time integration, inadequate support for multiple programming languages, insufficient collaboration features, and poor scalability for large codebases.
Modern software development practices are heavily reliant on continuous integration (CI) and continuous deployment (CD) pipelines 121, alongside integrated development environments (IDEs) 122 such as IntelliJ IDEA, Visual Studio Code, and Eclipse. However, most SATD detection tools lack seamless integration with these environments, necessitating manual execution or batch processing. This disrupts established workflows and reduces the tool’s utility for practitioners who require real-time feedback during development. The absence of real-time notifications limits the capacity of developers to address technical debt in a timely manner, thereby contributing to its accumulation. Tools designed for practical use should therefore provide IDE integration and support for automated scanning within CI/CD pipelines to enhance workflow efficiency and developer productivity.
Furthermore, existing tools often struggle to scale effectively in large and complex software projects. Many of these tools are designed with a narrow focus on specific programming languages, particularly Java, which limits their applicability in diverse development environments where teams work with multiple languages, such as Python, C++, and JavaScript. This restriction reduces their practicality for organizations seeking comprehensive SATD detection across their entire codebase. Performance optimization for large repositories and multi-language compatibility are thus critical areas for future improvement. In addition, many SATD detection tools rely on ML or DL models, which are often perceived as “black boxes”. This lack of interpretability may make it challenging for developers to fully trust and adopt these tools. Offering more transparent and explainable outputs—such as highlighted patterns, relevant keywords, or confidence scores—could enhance clarity, build trust, and encourage broader adoption among practitioners.
Moreover, existing tools seldom offer functionality for long-term monitoring and trend analysis of SATD. For projects with extended lifecycles, it is valuable to track how technical debt is introduced, accumulated, and eventually repaid. While tools such as DebtViz aim to visualize and monitor SATD over time, they remain in early stages of development and may not yet be suitable for large-scale applications. Enhancing these tools with features for historical analysis and prioritization of technical debt repayment could support more informed decision-making and strategic refactoring efforts.
These limitations highlight a broader challenge: the gap between academic research and industry needs. Many tools are developed with a strong focus on academic performance metrics, yet they may not fully address practical constraints, including integration, scalability, and usability. Strengthening collaboration between academia and industry is crucial to ensuring that SATD detection tools are not only scientifically rigorous but also practical and widely applicable in real-world software development settings.
5.4 Decision-making guidance for adopting SATD detection techniques
To address the lack of actionable guidance for adoption, this subsection synthesizes evidence from our review into practical decision considerations for selecting SATD detection techniques. Below, we summarize the findings based on the techniques, approaches, or models that were identified in our study. We also summarize the main advantages and disadvantages of each technique or approach.
  • Heuristic-based techniques (e.g., tagging rules, regular expressions, pattern matching) are most appropriate when teams need a lightweight solution with minimal infrastructure and strong interpretability. They are typically easy to implement, explain, and integrate into existing workflows, making them suitable for small-to-medium codebases or for quick triage (e.g., flagging obvious TODO/FIXME-like debt markers). However, the review evidence indicates that heuristics often miss more subtle SATD expressions and are sensitive to vocabulary differences across projects and developer writing styles. Therefore, heuristics are best positioned as a baseline or as a high-precision filter when false positives are costly, rather than as a comprehensive detector.
    - Advantages: Lightweight, easy to use
    - Disadvantages: Lower recall, limited coverage
  • Classic ML approaches (e.g., SVM, RF, and LR) offer a pragmatic middle ground for teams that have access to labeled data and seek improvements over heuristics without the full operational burden of deep models. Across the reviewed studies, ML methods tend to perform more robustly than heuristics while remaining easier to reproduce and debug than DL/Transformer approaches. ML is often a sensible choice for organizations that can maintain a modest labeling pipeline and require reasonably transparent features (e.g., TF-IDF-based models) to secure stakeholder buy-in. Nevertheless, reported performance remains highly dependent on dataset characteristics (e.g., class imbalance, artifact type, and preprocessing), and practitioners should expect re-tuning when moving to new domains.
    - Advantages: More robust than heuristics, can be reasonably interpretable (feature-based)
    - Disadvantages: Requires labeled data, performance may degrade under domain shift
  • Deep learning approaches (e.g., CNN/LSTM/GNN) are most appropriate when the goal is improved effectiveness under complex linguistic variation or when large training corpora are available. The review evidence shows that DL methods often improve detection quality relative to ML in within-project evaluations, but they entail higher operational costs (training time, hardware, maintenance) and reduced transparency. DL models are therefore best suited to teams that can sustain the MLOps overhead, have sufficient data volume, and accept that model explanations may require additional techniques (e.g., post hoc explainability) to support developers’ trust.
    - Advantages: Often higher effectiveness than classic ML, can leverage larger corpora and richer representations
    - Disadvantages: Higher compute and MLOps cost, lower transparency
  • Transformer-based models (e.g., BERT-family and related variants) are typically justified when SATD detection is a high-value, high-scale activity and teams prioritize effectiveness and broader language understanding. In the reviewed literature, Transformers often achieve the strongest reported results for SATD identification and are increasingly used across diverse textual artifacts. The main practical constraints are operational: they require more computational resources, more specialized expertise, and careful engineering for deployment and monitoring. In addition, their “black-box” nature can hinder adoption unless paired with interpretability and documentation practices. Practitioners should treat Transformers as a strategic investment: most suitable for large projects, multi-artifact pipelines, or scenarios in which missing SATD instances (false negatives) incur high downstream costs.
    - Advantages: State-of-the-art effectiveness, strong language understanding across artifacts
    - Disadvantages: Highest compute and expertise requirements, interpretability and deployment complexity
These qualitative observations are consistent with the overall performance patterns synthesized in Section 4.3, where more recent DL and Transformer-based methods tend to report higher precision/recall/F1 than earlier heuristic and classical ML approaches, albeit under heterogeneous datasets and evaluation protocols.
Our synthesis indicates that no single technique dominates across all criteria; the best choice depends on project constraints (e.g., available labeled data, compute budget, and expertise) and the intended use of the detector (e.g., quick triage vs. high recall detection). Based on the cross-study synthesis, practitioners can use the following cues to narrow down their choice: if the goal is immediate, low-cost triage with maximum explainability, heuristic-based techniques are typically the best starting point. If labeled data are available and reproducibility and maintainability are important, classic ML provides a strong and practical baseline. If performance requirements are higher and sufficient data/compute resources exist, DL techniques become appropriate. Finally, if an organization can support the operational overhead and needs state-of-the-art effectiveness across heterogeneous artifacts, Transformer-based models are generally the most suitable option.
5.5 Implications for practitioners
BERT-based models demonstrate superior precision and recall in SATD detection: by adopting these models, practitioners can significantly reduce their reliance on manual review processes, which are often time-consuming, inconsistent, and prone to human error. Automated approaches streamline the identification of technical debt, ensuring that critical SATD instances are accurately detected without the extensive overhead of manual analysis. This enables software teams to allocate their resources more efficiently and to focus on higher-value activities, such as refactoring and feature development.
However, one of the key challenges in SATD identification is class imbalance, where SATD instances are significantly outnumbered by non-SATD instances. This imbalance often leads detection models to favor the dominant (non-SATD) class, resulting in missed detections. Additionally, class imbalance affects SATD categorization, as SATD must be classified into specific types of technical debt, such as design debt, architecture debt, requirements debt, and others. In both cases, the imbalance increases the risk of false negatives 117, where important SATD instances are overlooked. To mitigate this issue, practitioners should employ strategies such as data augmentation (e.g., by creating additional synthetic SATD samples to balance the dataset) and domain-adaptation techniques, which optimize model performance across diverse projects and environments. These approaches help ensure that models maintain high accuracy and recall, even when faced with the imbalanced data typical of real-world software projects.
As outlined in subsection 5.4, we present a comprehensive framework to assist practitioners in selecting the most appropriate SATD detection approach tailored to their specific project needs and constraints. This framework enables practitioners to effectively balance trade-offs involving scalability, explainability, and resource demands. These factors are essential for the successful adoption of SATD detection tools.
Additionally, automated SATD detection and early identification of technical debt contribute to the sustainability of software systems. Since SATD is prevalent across many software projects (affecting between 2.4% and 31% of files 8), it presents a long-term challenge. SATD often persists for extended periods, with a median lifespan ranging from 18 to 172 days and, in some cases, lasting through more than 1,000 commits 123. By adopting automated detection techniques, organizations that adopt these techniques will likely see measurable improvements in development efficiency, reduced technical debt accumulation, and enhanced product quality. These benefits help mitigate the risks associated with delayed debt repayment.
@@ 5.6 Future directions
To address the challenges identified in this review, future research should prioritize enhancing the generalizability, explainability, and efficiency of SATD detection models. Techniques such as transfer learning, which allow models to be trained on one dataset and then applied to others with minimal retraining, could significantly improve generalizability. At the same time, incorporating explainable AI (XAI) techniques is crucial to ensure that the decision-making processes of these models are transparent and understandable for developers. Developing more interpretable models or employing post-hoc explanation techniques could foster trust in these tools and encourage their adoption in real-world projects. Future work should focus on improving the interpretability of deep learning and Transformer models, for example, by using visualizations, highlighting informative patterns, and providing simplified intermediate representations. These improvements would help practitioners better understand the models, build trust in their outputs, and use them more effectively in software development contexts.
Beyond improving models, future research should prioritize SATD detection tools that are scalable, easy to use, and designed for real-world adoption. In practice, this means tools that integrate smoothly with existing development environments, support team collaboration, and work across diverse software ecosystems, including multi-language codebases. Since modern workflows are centered on IDEs and CI/CD pipelines, SATD detectors should be deployable in both settings and provide timely, actionable signals. In particular, real-time or near-real-time feedback can help developers identify and address technical debt as part of their normal workflow, rather than as a follow-up activity performed after implementation. By closing the gap between research prototypes and production-ready tooling, the software engineering community can ensure that SATD detection techniques deliver tangible value for long-term software maintenance and evolution.
While our findings provide valuable insights into SATD detection, it is essential to acknowledge potential limitations that may influence the generalizability of our conclusions. The next section discusses threats to validity, including biases in study selection, dataset heterogeneity, and variations in evaluation metrics, and how we have addressed these concerns.
@@ 6 Threats to Validity
As with any systematic literature review, the results of this study are subject to various validity threats. These are categorized as internal, external, and construct validity threats, along with mitigation strategies to minimize their impact.
@@ 6.1 Internal validity
Internal validity refers to the reliability of the methodology employed in the study. One key threat to internal validity in this SLR is related to the formulation of the search strings. In the pilot search, several iterations of search strings were tested, starting with a more general query (e.g., “self-admitted technical debt” OR SATD and NLP), which yielded a broader, less focused set of results. Over time, more refined strings were developed, including terms such as “detection” and “identification”, as well as the use of Boolean operators and the wildcard (*) for broader term variations.
The final search string (S3), which combined terms such as “self-admitted technical debt” OR SATD with NLP OR “natural language processing” AND (detection OR identification OR prediction), was chosen after multiple pilot tests to ensure it captured the most relevant studies. This iterative process helped mitigate the risk of missing key studies during the initial literature search.
@@ 6.2 External validity
External validity refers to the extent to which the findings of this SLR can be generalized to other contexts beyond the studies reviewed. A primary threat to external validity in this research is the limited generalizability of the results due to the scope of the reviewed datasets. All studies included in this review rely on open-source projects, which may not fully capture the complexities or constraints encountered in proprietary or large-scale industrial environments.
Open-source datasets often have characteristics, such as well-documented codebases or community-driven development practices, that differ significantly from proprietary software projects, which may involve stricter deadlines, unique domain requirements, or higher security and compliance constraints. As a result, some of the trends and findings related to SATD detection may not extend to these more complex and regulated environments. For example, the performance of ML or DL models, which may perform well on open-source datasets, could degrade when applied to more diverse or proprietary datasets with varying development practices.
Another potential limitation of this review is that the synthesized results for RQ3 are based exclusively on within-project evaluations (e.g., k-fold cross-validation or random splits on the same project). While this choice ensures comparability across studies, it may overestimate performance relative to cross-project or out-of-distribution evaluations, which better reflect real-world deployment scenarios. Some primary studies did report cross-project results, but these were not included in our synthesis. Therefore, the performance metrics presented in this review should be interpreted with caution, as they represent results under controlled within-project settings rather than true generalizability across projects.
@@ 6.3 Construct validity
Construct validity concerns whether a study accurately measures the concepts it aims to investigate. A significant threat in this SLR is the comparison of results across studies, which can be misleading if studies employ different methodologies or report metrics inconsistently. Directly comparing the precision, recall, and F1 scores of models trained on heterogeneous datasets without considering their context may lead to inaccurate conclusions.
To address this, we employed a grouping strategy to compare similar studies and ensured that the SATD tasks were organized before any comparisons were made. By contextualizing the results within groups, we aimed to make the comparisons more meaningful. Furthermore, this methodology is based on established practices within the field, exemplified by Hall et al. 112 and Malhotra 113, which similarly compare results across studies to derive broad-based insights into algorithmic performance across varying contexts.
Having considered the potential limitations of this study, we now summarize our key findings and contributions. The final section revisits the research questions, highlights the main takeaways, and outlines directions for future work to advance SATD detection research and tool development.